080513 Al, Bob 0524 N test clutter scans: .24, .5, .75. 1.0 @ 3 deg/sec 0545 Y scans at 1, .2, .5, .75, 1.0 @ 3 deg/sec 0554 N (CF "off" part way thru previous 1.0 deg scan) Variable CF start nos. 0622 Y RHIs show some waves at the top of a 1.2KM B.L. 0830 Box Solar
080514 Mike, Al, Brian 0015 Started Box Solars 0139 Started transmitter for Tx power measure schek. Ran clutter scan without filters. 0200 Clutter Scan with clutter filters 0223 Y Running Sequence 5 0511 N Refractivity sweep at .3 and .4 deg elevs, 8 deb/sec (set reqd for 110 hits/beam, display shows 106) 0516 N Modified tilts to include .3, .4, .5 (run for 2hrs)
080515 Gales, Strong 0057 Y Test scans on sequence 5
080517 Gales 0615 Y Been testing scan controler for missing elev angles. 0656 Y Stopped scanning, disabled transmitter. Ran solar (unsuccessfully) 0710 Y Started transmitter and scan.
080518 Dixon, Strong 00:20 - 00:50 Solar Scan 0005 Arrived at tadar, default was running. note some sea clutter echo 0010 N Shut down transmitter, in preparation for solar. Solar did not work using the TIMREX - preIOP antenna file. 0020 N Changed to use TIMREX-1 scan file. Started solar. Seems to work OK. 0050 N Stopped Sun Scans 0052 N Started Simultaneous mode transmission test. Note: Second trip to SE - recorded time series 0116 N Stopped Simultaneous transmission - recorded time series 0120 Y Started normal ops. Start time series recording - low level "su--lace" 0143 Y Stop scanning. Stop time series recording. 0150 Y Testing scan using antenna controller. 0305 Changed VIRAQ Zdr fudge factor to -0.25
080519 Gales 23:46 - 23:58 Light rain 01:02 - 01:24 Solar Scan 2330 Y Arrived on site. Radar running unattended.(18 May UTC) 0102 N Disabled transmitter. Started Solar, ran OK. 0118 Changed VIRAQ Adr fudge factor to 0. 0124 N Started transmitter. Ran TIMREX-nonIOP scn. sequence 7 0130 Y Started clutter filter. Non-IOP 1630 Y Unattended Sequence
080520 Strong, Gales 05:50 - 05:58 Light rain 0132 Y Found clutter filter off (possibly turned off during morning's vertical pointing or software testing). Turned clutter filter on. Weather - light, scattered precip. 1420 Y TIMREX non-IOP Unattended SUR - occasional rain - light
080521 Gales, Strong 18:53 18:53 Light rain 0018 N VIRAQ locked, Hyperterminal window unresponsive. Stoped and restarted 0056 Antenna Stowed, rebooting VIRAQ PC. 0105 Y VIRAQ rebooted. Radar operational again.
080522 Strong, Gordon 0000 Y Unattended Sequences ? Y Added 16 degree tilt to unattended scan sequence 0737 Y Changed to ssite_all scan sequence 0754 Y Change scan sequence to RHIs at 15 deg azimuth to perform coordinated scans with TEAM Radar 0800 Y Changed back to ssite_all 0915 Y Changed to RHI - 45d. 1040 Y Unattended 1040 Y Sec-180 (40 deg - 90deg) 1100 Y Unattended 0945 Y RHI dz 70 deb ele 0 - 90deg with TEAM-R 1015 Y 37deb ele 0 - 60deb with TEAM-R 1230 Y Original Unattended Sequence.
080523 Gales,Gordon,Strong 2345 S-Pol down for routine maintenance (check the pedestal level and belt tensions) and to replace PMAC card to try to fix volume and scan number problem (dual part ran bit 2 seems to be stuck on spare PMAC card). 0724 Y S-Pol UP! 0750 Y running RHI-45 (PPI:35deg - 80deg 15ele, RHI 40deb-60deg) 1017 Volume numbers appear to be sequential.
080524 Gordon 0132 Y S-Pol operated unattended throughout the night. No apparent problems.
080525 Gales 0000 Pedestal Maint. 0012 Stop RDAs. Test ATE monitoring program 0248 Y S-Pol returned to operation. ATE monitoring program running.
080526 Gales,Strong 0000 Y Pedestal Maint. Return to unattended
080527 Gales 23:32 - 23:42 Solar Scans 0001 Y New IOP started on VIRAQ 2345 Y New IOP started on VIRAQ 2355 Reset rain accumulation ("killall rain" on pgen)
080528 Gales 03:30 - 03:35 Light rain
080529 Strong,Gales 08:49 - 008:58 Light rain 0023 Y S Site_all 0445 S-Pol down to swap Mitch Switch 0615 Y S-Pol up: operating with spare Mitch switch 0849 Vertical poointing 0858 Y Running unattended
080530 Mike,Gordon 0000 Y Rain accumulations are now reset at 00:00 Taiwan Standard Time (16:00 UTC) every day.
080531 Strong,Gales 09:23 - 09:29 Light rain 09:43 - 09:52 Medium rain 0000 Y Continue EOP2 1515 Y Light, scattered precip. 2333 Changed VIRAQ Zdr bias to -0.23 (config.rdr) 2336 Y Restored VIRAQ operation
080601 Strong, Gales 16:15 - 16:21 Light rain 0000 Y Continue EOP2 1615 Good vertical pointing data (I think)
080602 Strong, Gales 18:30 - 18:36 Light rain 0000 Y Coninue EOP2 0217 Changed TP phase setting 0210 Noticed Vert test puls missing ~0100 - 0218 0500 Mike Dixon (change/fix) to iagl covariance scaling for netCDF sweepfiles 0607 STOP (HAWK) radar for simultaneous HV scans 0611 Simultaneous HV mode 0614 Restart HAWK for alt. HV 0617 Alt HV mode 1838 Y Missed synchronization. Running default SUR scan 1845 Y Running unattended sequence
080603 Jonathan, Kyle 21:15 - 21:23 Light/Medium rain
080604 Jonathan, Kyle 06:23 - 06:36 Heavy rain 14:28 - 14:37 Medium rain 0216 Y Change RDA tapes
080605 Jonathan, Kyle 0458 Start Iq2Dsr.sim (simultaneous HV tx) 0459 Operating in SHV 0509 Restore to alternating HV mode 0510 S-Pol down 1448 Y S-Pol restored to operation
080606 Jonathan,Gordon 00:45 - 00:55 Solar Scans 09:21 - 09:35 Solar Scans 0056 Y S-Pol operating again after solar scans 0100 Y Running "unattended" scanning strategy 0921 N Solar Scans 0938 Y Running "unattended" scanning strategy
080607 Kyle, Tolly,Gordon 0113 Stow Ant for Maint. 0149 Y S-Pol operating 0308 Installed VIRAQ Info program on Mgen (replaces sweep-info broadcast, etc.) 2324 Running modified "refract" scan under cal-scans.scn for Rita. Only 0.5 deg elev 2350 Turned off transmitter pw RF to test tx power info in data stream
080608 Gordon,Kyle,Jonathan 0003 Y Turned on the RF input. 0034 Y Stopped REFRACT scan 0036 Y Start REFRACT scan 0055 Y Start modified REFRACT scan to work around stupid PMAC program bug (PMAC programs don't seem to turn off transition flag when repeating fixed angle - FIX) 0106 N Turn clutter filter off for refractivity scans. 0202 Y Stop refractivity scans 2330 N Running "REFRACT" 0.5deg and 1.1deg el for target ID
080609 Jonathan,Kyle,Gordon 0143 Y Stopped "REFRACT" target ID, Started TiMREX-nonIOP, unattended
080610 Gordon 0734 Y Continue operating - no IOP
080611 Jonathan,Gordon,Kyle 0635 Y No IOP - operating in the unattended scanning strategy
080612 Jonathan, Gordon, Kyle 0642 Camera - TV crew turned off SCC UPS. Sudden loss of all UPS Power 0642 Ejected RDA tapes manually. P25930, Q25930 0724 New IOP, started RDA's
080613 Gordon, Kyle, Jonathan - no log entry -
080614 Gordon, Kyle, Jonathan 12:09 - 12:12 Light rain - no log entry -
080615 Jonathan, Kyle, Gordon 01:21 - 01:25 Light rain 0353 HAWK stopped recording data (FM! probblem) 0502 HAWK restored to operation. (moments FMQ file removed, Iq2Dsr restarted) 0513 Switch to simultaneous HV mode 0517 Start sin HV time series recording 0538 Switch to alternating HV mode 0540 Operating in alt. HV mode and time series recording. 0546 Started new version of Iq2DS! (hopefully this will fix the FMQ problems) 0600 Stopped Time Series 0624 System Restored (WG switches in wrong position) 0642 Stopped time series recording ~1900 Changed pid tables to limit spurious grappel IDs
080616 Jonathan, Gordon, Kyle 0043 Start time series recording in preparation 0050 Start SHV recording 0051 Transmitter switch to SHV 0059 Y Restored operation to alt. HV
080617 Jonathan, Gordon 0440 El brake tripped. sweepfiles between 0440 - 0511 UTC are listed with a fixed angle of 3.6deb, but actually recorded with antenna angle of 2.6deg 0511 Restored to "unattended"
080618 Gordon, Kyle, Jonathan 00:45 - 00:59 Solar Scans 01:07 - 01:12 Solar Scans 02:10 - 02:40 ATE Calibration 0113 SPOL Down for maintenance 0247 SPOL back up ~2130 Mike Dixon intalled PID algorithm (0619?) 0718 modified legacy PID to relax 0degC thresholds
080619 Jonathan,Kyle,Gordon - no log entry -
080620 Kyle, Jonathan, Gordon - no log entry -
080621 Kyle, Gordon,Jonathan 0745 Refractt cal scans start 0907 Refractt cal scans stop
080622 Kyle, Jonathan,Gordon - no log entry -
080623 Kyle,Jonathan,Gordon - no log entry -
080624 Kyle,Jonathan,Gordon 0515 Refractt scans start 0615 Refractt scans stop
080625 Kyle,Jonathan,Gordon 0028 began using new refractivity Cal scan 1401 Modified threshold in refractivity input for stricter tolerances
080626 Kyle,Gordon - no log entry -
080627 Kyle,Jonathan - no log entry -
080628 Gordon *07:29 - 07:35 Rain 07:35 - 07:41 Rain 0721 Y Switch to simultaneous HV *0729 Vertical pointing in SHV 0735 Vertical pointing in alt. HV (Switched back to alternating HV during vertical pointing) 0743
080629 Kyle,Jonathan - no log entry -
080630 Kyle,Jonathan,Gordon 00:51 - 01:07 Solar Scans 0051 Sunbox solar 0051 -> 0107 0212 Finish ATE cal.
Special thanks to all the technicians, operators, and scientists for writing relevant information in this log. Even the Engineers wrote some comments.
Those participating through log entries are:Jonathan Emmet, Gordon Farquharson, Kyle Holden, Mike Strong, and Bryan Gales
This log was transcribed by Jean Hurst, and edited by Bob Rilling.
All but minor editorial changes to the log are indicated within square brackets.